# REINFORCE++: An Efficient RLHF Algorithm with Robustness to Both Prompt and Reward Models

### v5

@@ 5 Conclusion
This paper presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
Appendix A Algorithm Details
REINFORCE++-baseline.
For ease of use and reproducibility, we implement our proposed method within the OpenRLHF framework (https://github.com/OpenRLHF/OpenRLHF).
Our experiments above showed that training models using multiple generated responses per prompt can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards. The advantage for each prompt-response pair is then defined as:
$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$
$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}-\text{mean}_{batch}(Adv_{q,o_{t}})}{\text{std}_{batch}(Adv_{q,o_{t}})}$
where $group$ denotes the generated responses corresponding to the same prompt. This baseline calculation is similar to the approach taken by GRPO; however, we deliberately move the standard deviation ($std$) of the group normalization to the global batch normalization. In practice, we observed that including the standard deviation introduces training instability and exacerbates overfitting to easier prompts in the RLHF process. For the KL constraint, REINFORCE++-baseline uses $k_{2}$ instead of the $k_{3}$ divergence approximation method in GRPO. For more details, please refer to the next paragraph.
KL Penalty Design.
In the design of REINFORCE-based RLHF methods, the choice of KL divergence estimation significantly influences algorithmic stability and accuracy. The widely utilized $k_{3}$ estimation in GRPO, defined by the loss function:
@@ Appendix A Algorithm Details
$\mathcal{L}_{k_{3}}=\mathbb{E}_{x\sim D,y\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left(\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-\log\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-1\right)$ (12)
Which involve inherent biases due to its approximation nature. Specifically, the gradient of $k_{3}$ estimation approximates the gradient of the $k_{2}$ estimation through linearization:
@@
where $x=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$. Although the $k_{3}$ gradient aligns closely with the $k_{2}$ gradient when the policy is close to the reference ($x\approx 1$), substantial deviations result in two key deficiencies:
  1. 1.
    Bias: As the current policy diverges significantly from the reference, particularly when $\pi_{\text{ref}}\gg\pi_{\theta_{\text{old}}}$, the linear approximation error expands nonlinearly, causing considerable biases.
  2. 2.
    Asymmetry: The function $(x-1)$ exhibits asymmetric behavior depending on whether $\pi_{\theta_{\text{old}}}$ surpasses or falls below $\pi_{\text{ref}}$, further complicating convergence properties.
Therefore, despite its computational simplicity, $k_{3}$ estimation introduces biases and increased variance, suggesting it is not strictly superior to the theoretically unbiased $k_{2}$estimation. Experimental evidence consistently reveals that employing $k_{3}$ estimation in GRPO results in higher variance fluctuations than $k_{2}$ estimation. Further discussion can be found in Liu (2025).
Relationship with PPO.
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term. At the same time, the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
Other Implementation Details
@@
Appendix B Acknowledgements
  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
@@ Appendix B Acknowledgements
  • Wei Shen: Led the paper writing and supervised the main experiments.

### v8

@@ 6 Conclusion
In this paper, we presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global advantage normalization to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, and extending REINFORCE++ beyond RLHF settings.
Appendix A Proof: The GRPO Advantage Estimator is Biased
A.1 Assumptions and Settings
We observe $N$ rewards $r_{i}$ for a prompt, and assume the true baseline is $\theta$, such that
$r_{i}=\theta+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),\quad i=1,\dots,N,$
with all advantage value $\epsilon_{i}$ independent. Define
$\bar{\epsilon}=\frac{1}{N}\sum_{j=1}^{N}\epsilon_{j},\quad D=\sqrt{\frac{1}{N}\sum_{j=1}^{N}(\epsilon_{j}-\bar{\epsilon})^{2}},\quad A_{i}=\frac{\epsilon_{i}-\bar{\epsilon}}{D}.$
We will prove the following:
Theorem 1.
For any finite $N\geq 2$, the advantage estimator $A_{i}$ is biased:
$\mathbb{E}[A_{i}\mid\epsilon_{i}]\neq\epsilon_{i}.$
Proof.
Step 1: Numerator Bias
We rewrite the numerator:
$\epsilon_{i}-\bar{\epsilon}=\left(1-\frac{1}{N}\right)\epsilon_{i}-\frac{1}{N}\sum_{j\neq i}\epsilon_{j}.$
Since the $\epsilon_{j}$ for $j\neq i$ are zero-mean and independent of $\epsilon_{i}$,
$\mathbb{E}[\epsilon_{i}-\bar{\epsilon}\mid\epsilon_{i}]=\left(1-\frac{1}{N}\right)\epsilon_{i}.$
Step 2: Denominator Depends on $\epsilon_{i}$
(a) Compute $\mathbb{E}[D^{2}\mid\epsilon_{i}]$:
By definition,
$D^{2}=\frac{1}{N}\sum_{j=1}^{N}(\epsilon_{j}-\bar{\epsilon})^{2}=\frac{1}{N}\sum_{j=1}^{N}\epsilon_{j}^{2}-\bar{\epsilon}^{2}.$
Since
$\bar{\epsilon}=\frac{1}{N}\left(\epsilon_{i}+\sum_{j\neq i}\epsilon_{j}\right),$
and conditioning on $\epsilon_{i}$ keeps the $\epsilon_{j}$, $j\neq i$, i.i.d. $\mathcal{N}(0,\sigma^{2})$, we obtain:
$\mathbb{E}\left[\sum_{j=1}^{N}\epsilon_{j}^{2}\mid\epsilon_{i}\right]=\epsilon_{i}^{2}+(N-1)\sigma^{2},$
$\mathbb{E}[\bar{\epsilon}^{2}\mid\epsilon_{i}]$ $=\frac{1}{N^{2}}\mathbb{E}\left[\left(\epsilon_{i}+\sum_{j\neq i}\epsilon_{j}\right)^{2}\Big{|}\epsilon_{i}\right]$
$=\frac{1}{N^{2}}\left(\epsilon_{i}^{2}+2\epsilon_{i}\cdot\underbrace{\mathbb{E}\left[\sum_{j\neq i}\epsilon_{j}\right]}_{0}+\mathbb{E}\left[\left(\sum_{j\neq i}\epsilon_{j}\right)^{2}\right]\right)$
$=\frac{\epsilon_{i}^{2}+(N-1)\sigma^{2}}{N^{2}}.$
Subtracting:
$\mathbb{E}[D^{2}\mid\epsilon_{i}]$ $=\frac{1}{N}(\epsilon_{i}^{2}+(N-1)\sigma^{2})-\frac{\epsilon_{i}^{2}+(N-1)\sigma^{2}}{N^{2}}$
$=\underbrace{\frac{(N-1)^{2}}{N^{2}}\sigma^{2}}_{\alpha}+\underbrace{\frac{N-1}{N^{2}}}_{\beta}\epsilon_{i}^{2}=\alpha+\beta\epsilon_{i}^{2}.$ (15)
(b) $g(\epsilon_{i})$ is Not Constant:
Let $g(\epsilon_{i})=\mathbb{E}\left[1/D\mid\epsilon_{i}\right]$ and $\mu(\epsilon_{i})=\mathbb{E}[D^{2}\mid\epsilon_{i}]=\alpha+\beta\epsilon_{i}^{2}$.
Using Taylor expansion of $f(x)=1/\sqrt{x}$ around $x_{0}=\mu(\epsilon_{i})$:
$f(x)=\frac{1}{\sqrt{x_{0}}}-\frac{1}{2}\frac{x-x_{0}}{x_{0}^{3/2}}+\frac{3}{8}\frac{(x-x_{0})^{2}}{x_{0}^{5/2}}+O((x-x_{0})^{3})$
Taking conditional expectation:
@@ A.1 Assumptions and Settings
$g(\epsilon_{i})$ $=\mathbb{E}[1/D\mid\epsilon_{i}]=\mathbb{E}\left[\frac{1}{\sqrt{D^{2}}}\mid\epsilon_{i}\right]=\mathbb{E}\left[f(D^{2})\mid\epsilon_{i}\right],\quad\text{where }f(x)=\frac{1}{\sqrt{x}}$
$\approx f(\mu(\epsilon_{i}))+f^{\prime}(\mu(\epsilon_{i}))\cdot\mathbb{E}[D^{2}-\mu(\epsilon_{i})\mid\epsilon_{i}]+\frac{f^{\prime\prime}(\mu(\epsilon_{i}))}{2}\cdot\mathbb{E}[(D^{2}-\mu(\epsilon_{i}))^{2}\mid\epsilon_{i}]$
$=\frac{1}{\sqrt{\mu(\epsilon_{i})}}-\frac{1}{2\mu(\epsilon_{i})^{3/2}}\cdot\underbrace{\mathbb{E}[D^{2}-\mu(\epsilon_{i})\mid\epsilon_{i}]}_{=\,0}+\frac{3}{8\mu(\epsilon_{i})^{5/2}}\cdot\text{Var}(D^{2}\mid\epsilon_{i})$
$=\frac{1}{\sqrt{\mu(\epsilon_{i})}}+\frac{3}{8}\cdot\frac{\text{Var}(D^{2}\mid\epsilon_{i})}{\mu(\epsilon_{i})^{5/2}}$
Since $\mu(\epsilon_{i})=\alpha+\beta\epsilon_{i}^{2}$ with $\beta>0$, the first term alone shows that $g(\epsilon_{i})$ depends on $\epsilon_{i}^{2}$ and hence is not constant.
Step 3: Putting It Together
Decomposing $A_{i}$,
$A_{i}=\frac{\epsilon_{i}-\bar{\epsilon}}{D}=\left(1-\frac{1}{N}\right)\frac{\epsilon_{i}}{D}-\left(\frac{1}{N}\sum_{j\neq i}\epsilon_{j}\right)\cdot\frac{1}{D}.$
[…]
$\mathbb{E}\left[\left(\frac{-1}{N}\sum_{j\neq i}\epsilon_{j}\right)\cdot\frac{1}{D}\Big{|}\epsilon_{i}\right]=0.$
It follows that
$\mathbb{E}[A_{i}\mid\epsilon_{i}]=\left(1-\frac{1}{N}\right)\epsilon_{i}\cdot\mathbb{E}\left[\frac{1}{D}\mid\epsilon_{i}\right]=\left(1-\frac{1}{N}\right)\epsilon_{i}\cdot g(\epsilon_{i}).$
Step 4: Concluding the Bias
If $A_{i}$ were unbiased, we would have:
$\left(1-\frac{1}{N}\right)g(\epsilon_{i})\equiv 1\quad\Rightarrow\quad g(\epsilon_{i})\equiv\frac{N}{N-1},$
which contradicts Step 2. Therefore, for any finite $N\geq 2$,
$\mathbb{E}[A_{i}\mid\epsilon_{i}]\neq\epsilon_{i}.$
Hence $A_{i}$ is a biased estimator. ∎
A.2 Why Use Global Batch Normalization?
We observe that as $N\to\infty$, the denominator of the estimator converges to the constant $\sigma$, and the bias in the numerator vanishes. This insight underlies our decision to adopt global batch normalization in both REINFORCE++ and REINFORCE++-Baseline, as the global batch size is typically much larger than the group batch size in practice.
Appendix B Algorithm Details
B.1 KL Penalty Design
In the design of REINFORCE-based RLHF methods, the choice of KL-divergence estimation significantly influences algorithmic stability and accuracy. We briefly introducing the widely utilized KL estimators,
$\mathcal{L}_{k_{2}}$ $=\mathbb{E}_{s\sim D,\,a\sim\pi_{\theta_{\text{old}}}(\cdot|s)}\left(\frac{1}{2}(-\log x)^{2}\right)$
$\mathcal{L}_{k_{3}}$ $=\mathbb{E}_{s\sim D,\,a\sim\pi_{\theta_{\text{old}}}(\cdot|s)}\bigl{(}(x-1)-\log x\bigr{)}\qquad$
$\text{where}\ x$ $=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$
The widely utilized $k_{3}$ estimator in GRPO involves inherent biases due to its approximation nature. First, we directly provide the gradients of k2 and k3 with respect to the policy:
@@ B.1 KL Penalty Design
It is not hard to see that under the vanilla policy gradient ($r\cdot\nabla_{\theta}\log\pi_{\theta}$), the gradient with respect to k2 is equivalent to the gradient when placing the logarithmic KL term in the reward $r$, therefore, k2 is unbiased. When performing a Taylor expansion in the policy neighborhood ($x\approx 1$), $\log x\approx x-1$, At this time, the k3 gradient forms a linear approximation of the k2 gradient. However, this approximation has two key flaws:
  • Bias: When the policy significantly deviates from the reference policy ($x$ is far from 1, especially in the later stages of training when $\pi_{\text{ref}}$ is far from $\pi_{\theta_{\text{old}}}$, particularly when $\pi_{\text{ref}}>>\pi_{\theta_{\text{old}}}$), the approximation error grows nonlinearly.
  • Asymmetry: The response characteristics of $x-1$ are asymmetric for $\pi_{\theta_{\text{old}}}>\pi_{\text{ref}}$ and $\pi_{\theta_{\text{old}}}<\pi_{\text{ref}}$.Therefore, "k3 estimation as a loss function" is merely an approximation of "k2 estimation as a loss function".
Therefore, despite its computational simplicity, $k_{3}$ estimation introduces biases and increased variance, suggesting it is not strictly superior to the theoretically unbiased $k_{2}$ estimation. Experimental evidence consistently reveals that employing $k_{3}$ estimation in GRPO results in higher variance fluctuations than $k_{2}$ estimation. Further discussion can be found in (Liu, 2025).
B.2 Implementation Tricks
@@
Appendix C Tool-Integrated Reasoning and Agent Experiments
In this section, we present a comprehensive evaluation of the effectiveness of various reinforcement learning algorithms in the context of tool-augmented reasoning. Our experiments are conducted within the framework established by Mai et al. (2025), which features a zero-shot agent environment designed for large language models to tackle mathematical problems. This environment allows the models to invoke external tools, such as Python execution, without the need for supervised demonstrations of tool usage.
To ensure a fair comparison, we adhere to the training and evaluation protocols of the ZeroTIR environment. The backbone model utilized in our experiments is Qwen 2.5 Base 7B, which is initialized from pre-trained weights to maintain consistency across trials. The training process is facilitated by OpenRLHF Hu et al. (2025a), utilizing datasets from ORZ Hu et al. (2025b) and DAPO Yu et al. (2025).
We assess the performance of the models on several established mathematical benchmarks, including AIME 2024 Jia , AIME 2025 math ai , HMMT FEB 2024/2025, and CMIMC. The evaluation metric employed is the average@32, consistent with the methodology outlined in Mai et al. (2025).
[…]
The results of our experiments, summarized in Table 5, indicate that the REINFORCE++-baseline algorithm consistently outperforms both GRPO and PPO across all benchmarks, achieving the highest average accuracy. This finding aligns with previous research by Mai et al. (2025), which highlighted the stability and effectiveness of the REINFORCE++-baseline and global advantage normalization in tool-augmented environments.
Algorithm AIME 24 AIME 25 HMMT 2025 HMMT 2024 CMIMC Avg
GRPO 31.66 21.87 16.97 17.70 24.68 22.58
PPO 30.20 21.66 15.00 18.43 23.95 21.85
RF++-baseline 30.83 27.18 17.91 18.95 25.62 24.10
Table 5: Performance Comparison Across Benchmarks (average@32).
Appendix D Best Practices
This paper introduces two algorithms: REINFORCE++ and REINFORCE++-Baseline, with a key focus on selecting the appropriate algorithm based on task conditions.
Empirical evidence from the open-source community suggests that REINFORCE++-Baseline is particularly effective for sample filtering or more complex scenarios. For instance, in the multi-turn tool-calling setting discussed earlier, a high proportion of void (i.e., non-informative) samples can destabilize training. In such cases, incorporating a baseline—by subtracting the intra-group mean reward—significantly improves training stability by effectively filtering out void samples. Moreover, automatic reward reshaping reduces the need for designing complex reward structures. For example, REINFORCE++-Baseline supports both 0/1 and -1/1 reward schemes, while REINFORCE++ performs best with symmetric rewards, such as -1/1 in RLVR tasks.
In contrast, for tasks where obtaining multiple reward signals for distinct responses conditioned on the same intermediate state is challenging—such as training with Process-Supervised Reward Models (PRMs) or online real-time sampling—we recommend using REINFORCE++.
Appendix E Acknowledgements
  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the proof of GRPO advantage estimator.
@@ Appendix E Acknowledgements
  • Haotian Xu: Conducted comparative experiments between reinforce++-baseline and complex tool calling as well as agent scenarios.
  • Wei Shen: Led the paper writing and designed and supervised the main experiments.